OpenStudy (liliakarina):

If x+3/2+x= 2/3, then x=? I say it's -1, but it doesn't work for 2+x

OpenStudy (helder_edwin):

is you equation \[ \Large \frac{x+3}{2+x}=\frac{2}{3} \]

OpenStudy (liliakarina):

Yes, exactly

OpenStudy (helder_edwin):